A random variable X has the following density function fx(x) = 2 − kx, 0 < x < 1/k 

a) Find the constant k.

b) Find the mean.

c) Find the standard deviation.


1
Expert's answer
2021-08-23T15:30:50-0400

(a)


"\displaystyle\int_{-\infin}^{\infin}f(x)dx=\displaystyle\int_{0}^{1/k}(2-kx)dx=[2x-\dfrac{kx^2}{2}]\begin{matrix}
 1/k \\
 0
\end{matrix}"

"=2/k-1/(2k)=\dfrac{3}{2k}=1=>k=\dfrac{3}{2}"

(b)


"E(X)=\displaystyle\int_{-\infin}^{\infin}xf(x)dx=\displaystyle\int_{0}^{2/3}(2x-\dfrac{3}{2}x^2)dx"

"=[x^2-\dfrac{x^3}{2}]\begin{matrix}
 2/3 \\
 0
\end{matrix}=\dfrac{8}{27}"

(c)


"E(X^2)=\displaystyle\int_{-\infin}^{\infin}x^2f(x)dx=\displaystyle\int_{0}^{2/3}(2x^2-\dfrac{3}{2}x^3)dx"

"=[\dfrac{2}{3}x^3-\dfrac{3}{8}x^4]\begin{matrix}
 2/3 \\
 0
\end{matrix}=\dfrac{10}{81}"

"Var(X)=\sigma^2=E(X^2)-(E(X))^2"

"=\dfrac{10}{81}-(\dfrac{8}{27})^2=\dfrac{26}{729}"

"\sigma=\sqrt{\sigma^2}=\dfrac{\sqrt{26}}{27}"